Mike's age is 3 times john's age. in 12 years, mike will be twice as old as john.how old are mike and john now?

Answer by checkley77(12844)

M=3X
J=X
3X+12=2(X+12)
3X+12=2X+24
3X-2X=24-12
X=12 AGE OF JOHN NOW.
3*12=36 AGE OF MIKE NOW.
PROOF:
3*12+12=2(12+12)
36+12=2*24
48=48